News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 1 March 2002 

During the last week, Chandra completed the observing schedule as planned.

Chandra completed the last eclipe of the 2002 spring season on February 24 with nominal power and thermal performance. On February 25 a real time procedure was carried out to return the electrical power system parameters to their default, medium eclipse, settings.

On February 26, a calibration of the Aspect Camera Assembly's dark current was carried out by means of stored command loads, and was monitored in real time. This was the first time the calibration was automated in the daily loads, a significant accomplishment that makes it possible to schedule ACA dark current calibrations more frequently than previously, when real time commanding was used.

A press release was issued on February 27 featuring a pulsating X-ray hot spot on Jupiter that challenges existing theories of Jupiter's X-ray emission. See http://chandra.harvard.edu/press/02_releases/press_022702.html

The observing schedule for the next week is shown below. It includes an observation of 4U 1728-34 coordinated with BeppoSAX.

------------------------------------------
RCW103                 ACIS-S       Mar 3
AX J1844-02            HRC-I
Radiation Belts
4U 1728-34             ACIS-S/HETG  Mar 4
3EG J1835+5918         ACIS-S
NGC 1265               ACIS-S       Mar 5
Q0134+329              ACIS-S       Mar 6
Radiation Belts
V 410 TAU              ACIS-S       Mar 7
0745+241               ACIS-S
SDSS091316+591921      ACIS-S
NGC 4136 X-1           ACIS-S
IC 3599                ACIS-S
LBQS 1208+1535         ACIS-S
NGC 3507               ACIS-S
LBQS 1205+1436         ACIS-S       Mar 8
Q1143-245              ACIS-S
1258-321               ACIS-S
ESO3060170-B           ACIS-I
ESO3060170-A           ACIS-I
Radiation Belts                     Mar 9
Puppis N Knot          ACIS-S
MACSJ0417.5-1154       ACIS-I       Mar 10
MACSJ0308.9+2645       ACIS-I
NRGB045                ACIS-I

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports